OpenStudy (anonymous):

How do you solve this proportion 2m - 1 / 3 = m + 1 / 2

OpenStudy (anonymous):

You would cross multiply. \[\frac{2m - 1}{3} = \frac{m + 1}{2} \implies 2(2m - 1) = 3(m + 1)\]Can you solve from here?

OpenStudy (anonymous):

Would you just use the distributive property

OpenStudy (anonymous):

Yes. You would distribute the 2 and 3 and then solve for m.

OpenStudy (anonymous):

I forgot but would you multiply the 2 x -1 or no

OpenStudy (anonymous):

Would it be m = 5

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

So 5 is right

OpenStudy (anonymous):

\[\huge \textbf{Yes, m = 5 :)}\]

OpenStudy (anonymous):

Ok cool thanks

OpenStudy (anonymous):

np :)